### v1

@@ 4.1 Divergence-conforming approximation
$\boldsymbol{\mathrm{H}}_{h}$ :=$:=\left\{\boldsymbol{v}_{h}\in\mathrm{H}(\mathop{\mathrm{div}}\nolimits,\Omega)\;:\;\boldsymbol{v}|_{K}\in[\mathbb{P}_{k+1}(K)]^{n},\,K\in{\mathcal{T}}_{h},{\ (\boldsymbol{v}_{h}\cdot\boldsymbol{n})|_{e\in\mathcal{E}_{\mathrm{in}}}=\widehat{u},\ (\boldsymbol{v}_{h}\cdot\boldsymbol{n})|_{e\in\mathcal{E}_{\mathrm{wall}}}=0}\right\},$
$\boldsymbol{\mathrm{H}}_{h,0}$ :=$:=\left\{\boldsymbol{v}_{h}\in\mathrm{H}(\mathop{\mathrm{div}}\nolimits,\Omega)\;:\;\boldsymbol{v}|_{K}\in[\mathbb{P}_{k+1}(K)]^{n},\,K\in{\mathcal{T}}_{h},{\ (\boldsymbol{v}_{h}\cdot\boldsymbol{n})|_{e\in\mathcal{E}_{\mathrm{D}}}=0}\right\},$
$\boldsymbol{\mathrm{H}}_{h,0}^{0}$ :=$:=\left\{\boldsymbol{v}_{h}\in\mathrm{H}(\mathop{\mathrm{div}}\nolimits,\Omega)\;:\;\boldsymbol{v}|_{K}\in[\mathbb{P}_{k+1}(K)]^{n},\,K\in{\mathcal{T}}_{h},\;\mathop{\mathrm{div}}\nolimits(\boldsymbol{v}_{h})=0\;\mbox{in}\,\Omega,{\ (\boldsymbol{v}_{h}\cdot\boldsymbol{n})|_{e\in\mathcal{E}_{\mathrm{D}}}=0}\right\},$
$Q_{h}$ :=$:=\left\{q_{h}\in L^{2}(\Omega)\;:\;q|_{K}\in\mathbb{P}_{k}(K),\,K\in{\mathcal{T}}_{h}\right\},$
$W_{h}$ :=$:=\left\{\xi_{h}\in{\mathrm{H}^{-1/2}(\Gamma_{\mathrm{m}})}\;:\;\xi|_{\overline{\Gamma_{j}}}\in\mathbb{P}_{k}(\overline{\Gamma_{j}}),\,j=1,...,n_{\mathcal{E}_{m}}\right\},$
$Z_{h}$ :=$:=\left\{\tau_{h}\in Z\cap C(\overline{\Omega})\;:\;\tau|_{K}\in\mathbb{P}_{k+1}(K),\,K\in{\mathcal{T}}_{h}\right\}.$
Here, $\mathbb{P}_{r}(\mathcal{O})$, for $r\geq 0$, denotes the space of piecewise polynomials of degree less than or equal to $r$ defined on the entity $\mathcal{O}$, and $\widehat{u}\in\mathbb{P}_{r+1}(\Gamma_{\mathrm{in}})$ is an interpolation of $\boldsymbol{u}_{\mathrm{in}}\cdot\boldsymbol{n}$. For the discrete space of the Lagrange multiplier, we consider a triangulation of $\Gamma_{\mathrm{m}}$ given by $\{\Gamma_{j}\}_{j=1}^{n_{\mathcal{E}_{m}}}$, where $n_{\mathcal{E}_{m}}$ corresponds to the number of facets in $\Gamma_{\mathrm{m}}$. Note that the discrete space for the velocity is nonconforming in $\boldsymbol{\mathrm{H}}$, and correspond to well-known divergence-conforming BDM elements family (denoted by $\mathbb{BDM}_{k+1}$) (see [7]). As the discrete velocity now lives in $\mathrm{H}(\mathop{\mathrm{div}}\nolimits,\Omega)$ and its normal trace is in $\mathrm{H}^{-1/2}(\partial\Omega)$, the pairings on $\Gamma_{\mathrm{m}}$ from (7),(9) suggest a discrete Lagrange multiplier space conforming with $\mathrm{H}^{1/2}(\Gamma_{\mathrm{m}})$ instead of $\mathrm{H}^{-1/2}(\Gamma_{\mathrm{m}})$. In that case, in (9) one should use $\mathcal{R}^{-1}_{1/2}(g(\theta))$ instead of $g(\theta)$ (where $\mathcal{R}_{1/2}$ denotes the Riesz map between $\mathrm{H}^{-1/2}(\Gamma_{\mathrm{m}})$ and its dual). However, we maintain $W_{h}$ defined as conforming with $\mathrm{H}^{-1/2}(\Gamma_{\mathrm{m}})$ as in the previous section. We bear in mind that the off-diagonal bilinear form (to be denoted $b_{h}(\cdot,\cdot)$ in (35) below) will be therefore slightly different, needing the Riesz representative of the Lagrange multiplier. The remaining spaces $Q_{h}$ and $Z_{h}$ are conforming in $Q$ and $Z$, respectively. With this choice of spaces, we introduce discontinuous versions of the bilinear forms $a(\cdot,\cdot),b(\cdot,\cdot)$ and the trilinear form $\widetilde{a}(\cdot;\cdot,\cdot)$. For the first, we follow the symmetric interior penalty form given by
$a_{h}(\boldsymbol{u},\boldsymbol{v}):=$ $\,\mu_{0}\int_{\Omega}\nabla_{h}\boldsymbol{u}:\nabla_{h}\boldsymbol{v}-\sum_{e\in\mathcal{E}_{I}\cup\mathcal{E}_{D}}\mu_{0}\int_{e}\{\!\!\{\nabla_{h}\boldsymbol{u}\}\!\!\}:\llbracket\boldsymbol{v}\otimes\boldsymbol{n}_{e}\rrbracket$ (34)
@@
where $\alpha_{0}>0$ is the stabilisation parameter. The broken gradient operator $\nabla_{h}$ is defined by $\nabla_{h}\boldsymbol{u}=\nabla(\boldsymbol{u}|_{K})$ for all $K\in{\mathcal{T}}_{h}$. For the off-diagonal bilinear form we use the same functional form as $b(\cdot,\cdot)$ but the spaces are different due to the different pairings discussed above
$b_{h}(\boldsymbol{v},(q,\xi)):=-\int_{\Omega}q\nabla\cdot\boldsymbol{v}+{\langle\boldsymbol{v}\cdot\boldsymbol{n},\mathcal{R}_{1/2}\xi\rangle}_{\Gamma_{m}}\qquad\forall\boldsymbol{v}\in\boldsymbol{\mathrm{H}}_{h},(q,\xi)\in Q_{h}\times W_{h}.$ (35)
@@
$\widetilde{a}_{h}(\boldsymbol{w};\boldsymbol{u},\boldsymbol{v})=\rho_{0}\int_{\Omega}(\boldsymbol{w}\cdot\nabla\boldsymbol{u})\cdot\boldsymbol{v}+\frac{\rho_{0}}{2}\sum_{e\in\mathcal{E}_{I}}\int_{e}(\boldsymbol{w}\cdot\boldsymbol{n}_{e}-|\boldsymbol{w}\cdot\boldsymbol{n}_{e}|)(\boldsymbol{u}^{+}-\boldsymbol{u})\cdot\boldsymbol{v},$ (36)
where $\boldsymbol{u}^{+}$ is the trace of $\boldsymbol{u}$ pointing in the upwind direction. If $\boldsymbol{w}\in\boldsymbol{\mathrm{H}}_{h,0}^{0}$, then the following property holds:
@@
$a_{h}(\boldsymbol{u}_{h},\boldsymbol{v}_{h})+\widetilde{a}_{h}(\boldsymbol{u}_{h};\boldsymbol{u}_{h},\boldsymbol{v}_{h})+b_{h}(\boldsymbol{v}_{h},(p_{h},\lambda_{h}))$ $=F(\boldsymbol{v}_{h}),$ (37)
@@
We recall that for a straight membrane such as the one depicted in Figure 2.1, the lifting
$\boldsymbol{\mathrm{H}}^{1}(\Omega)\rightarrow\mathrm{H}^{1/2}(\Gamma_{\mathrm{m}});\quad\boldsymbol{v}\mapsto\boldsymbol{v}|_{\Gamma_{\mathrm{m}}}\cdot\boldsymbol{n},$
holds (in the sense of the continuity of right inverse of the trace operator). Similarly, the corresponding discrete lifting over a mesh of $\Omega$ and $\Gamma_{\mathrm{m}}$ also holds.
@@
$\|\boldsymbol{v}\|_{1,h}^{2}:=\|\boldsymbol{v}\|_{0,\Omega}^{2}+\|\nabla_{h}\boldsymbol{v}\|_{0,\Omega}^{2}+\sum_{e\in\mathcal{E}_{I}\cup\mathcal{E}_{D}}h_{e}^{-1}\|\llbracket\boldsymbol{v}\otimes\boldsymbol{n}_{e}\rrbracket\|^{2}_{0,e}.$ (38)
@@
The following bounds hold true {linenomath}
$|\widetilde{a}(\boldsymbol{w};\boldsymbol{u},\boldsymbol{v})|$ $\lesssim\|\boldsymbol{w}\|_{1,h}\,\|\boldsymbol{u}\|_{1,h}\,\|\boldsymbol{v}\|_{1,h},$ $\boldsymbol{w},\boldsymbol{u}\in\boldsymbol{\mathrm{H}}^{1}(\mathcal{T}_{h}),\boldsymbol{v}\in\boldsymbol{\mathrm{H}}_{h},$ (39a)
$|\widetilde{c}(\boldsymbol{w};\theta,\tau)|$ $\lesssim\|\boldsymbol{w}\|_{1,h}\,\|\tau\|_{1,\Omega}\,\,\|\theta\|_{1,\Omega},$ $\boldsymbol{w}\in\boldsymbol{\mathrm{H}}^{1}(\mathcal{T}_{h}),\theta,\tau\in Z_{h},$ (39b)
@@
$|\widetilde{a}(\boldsymbol{w};\boldsymbol{u},\boldsymbol{v})|$ $\lesssim\|\boldsymbol{w}\|_{\mathbf{L}^{4}(\Omega)}\,\|\boldsymbol{u}\|_{1,h}\,\|\boldsymbol{v}\|_{\mathbf{L}^{4}(\Omega)}$
$\leq\|\boldsymbol{w}\|_{1,h}\,\|\boldsymbol{u}\|_{1,h}\,\|\boldsymbol{v}\|_{1,h}.$
@@ 5 Well-posedness of the divergence-conforming discrete problem
[…]
@@
The first bound directly follows from [22, Prop. 10]. Using (13b) gives the second estimate.
@@
[…]
where
@@
The first bound directly follows from [22, Prop. 10]. The proof of the second inf-sup condition can be done similarly to [27, Corollary 3.5].
@@
In the following result we prove an inf-sup condition of the linear part of (37) that will be useful to ensure the uniqueness and convergence of the discrete solution.
@@
$|||(\boldsymbol{v},q,\tau,\xi)|||\lesssim|||(\boldsymbol{u}_{h},p_{h},\theta_{h},\lambda_{h})|||,$
@@
$|||(\boldsymbol{u}_{h},p_{h},\theta_{h},\lambda_{h})|||^{2}\lesssim B((\boldsymbol{u}_{h},p_{h},\theta_{h},\lambda_{h}),(\boldsymbol{v},q,\tau,\xi)),$ (40)
@@
${B(\boldsymbol{u}_{h},p_{h},\theta_{h},\lambda_{h};\boldsymbol{v},q,\tau,\xi)=a_{h}(\boldsymbol{u}_{h},\boldsymbol{v})+b_{h}(\boldsymbol{v},(p_{h},\lambda_{h}))+b_{h}(\boldsymbol{u}_{h},(q,\xi))+c(\theta_{h},\tau),}$
@@
$|||(\boldsymbol{v},q,\tau,\xi)|||^{2}:=\|\boldsymbol{v}\|_{1,h}^{2}+||q||_{0,\Omega}^{2}+\|\xi\|_{-1/2,\Gamma_{m,h}}^{2}+||\tau||_{1,\Omega}^{2}.$
@@
Let $(\boldsymbol{u}_{h},p_{h},\theta_{h},\lambda_{h})$ be a solution of (37). If $\|\boldsymbol{u}_{h}\|_{1,h}\leq M$, for sufficiently small positive $M<1$, then $(\boldsymbol{u}_{h},p_{h},\theta_{h},\lambda_{h})$ is the unique solution of (37). Moreover, the following estimate holds:
$|||(\boldsymbol{u}_{h},p_{h},\lambda_{h},\theta_{h})|||\lesssim g_{2}.$
@@
Let $(\boldsymbol{u}_{1},p_{1},\theta_{1},\lambda_{1})$ and $(u_{2},p_{2},\theta_{2},\lambda_{2})$ be two discrete weak solutions of 37. Using Lemma 5.7, for each $({\boldsymbol{u}}_{1}-\boldsymbol{u}_{2},{p}_{1}-p_{2},{\theta}_{1}-\theta_{2},{\lambda}_{1}-\lambda_{2})\in{\boldsymbol{\mathrm{H}}_{h,0}}\times Q_{h}\times Z_{h}\times W_{h}$, we find $(\boldsymbol{v},q,\tau,\xi)\in$ with
$|||(\boldsymbol{v},q,\tau,\xi)|||\lesssim|||({\boldsymbol{u}}_{1}-\boldsymbol{u}_{2},{p}_{1}-p_{2},{\theta}_{1}-\theta_{2},{\lambda}_{1}-\lambda_{2})|||,$
@@
$|||({\boldsymbol{u}}_{1}-\boldsymbol{u}_{2},{p}_{1}-p_{2},{\theta}_{1}-\theta_{2},{\lambda}_{1}-\lambda_{2})|||^{2}\lesssim B({\boldsymbol{u}}_{1}-\boldsymbol{u}_{2},{p}_{1}-p_{2},{\theta}_{1}-\theta_{2},{\lambda}_{1}-\lambda_{2};\boldsymbol{v},q,\tau,\xi),$
where
${B(\boldsymbol{u}_{h},p_{h},\theta_{h},\lambda_{h};\boldsymbol{v},q,\tau,\xi)=a_{h}(\boldsymbol{u}_{h},\boldsymbol{v})+b_{h}(\boldsymbol{v},(p_{h},\lambda_{h}))+b_{h}(\boldsymbol{u}_{h},(q,\xi))+c(\theta_{h},\tau)}.$
By (37), it follows:
$|||({\boldsymbol{u}}_{1}-\boldsymbol{u}_{2},{p}_{1}-p_{2},{\theta}_{1}-\theta_{2},{\lambda}_{1}-\lambda_{2})$ $|||^{2}\lesssim B({\boldsymbol{u}}_{1},{p}_{1},{\theta}_{1},{\lambda}_{1};\boldsymbol{v},q,\tau,\xi)-B(\boldsymbol{u}_{2},p_{2},\theta_{2},\lambda_{2};\boldsymbol{v},q,\tau,\xi)$
@@
$\quad+|\langle\xi,g(\theta_{1})-g(\theta_{2})\rangle_{\Gamma_{m}}|.$ (41)
@@
$|\widetilde{a}_{h}(\boldsymbol{u}_{1},\boldsymbol{u}_{1},\boldsymbol{v})-\widetilde{a}_{h}(\boldsymbol{u}_{2},\boldsymbol{u}_{2},\boldsymbol{v})|$ $\leq M\|(\boldsymbol{u}_{1}-{\boldsymbol{u}}_{2})\|_{1,h}\;\|\boldsymbol{v}\|_{1,h},$ (42a)
$|\widetilde{c}(\boldsymbol{u},\theta,\tau)-\widetilde{c}(\boldsymbol{u},\theta,\tau)|$ $\leq M\big{(}||\theta_{1}-{\theta}_{2}||_{1,\Omega}+\|(\boldsymbol{u}_{1}-{\boldsymbol{u}}_{2})\|_{1,h}\big{)}||\tau||_{1,\Omega},$ (42b)
$|\langle\xi,g(\theta_{1})-g(\theta_{2})\rangle_{\Gamma_{m}}|$ $\leq L||\xi||_{-1/2,\Gamma_{m,h}}||(\theta_{1}-{\theta}_{2})||_{1,\Omega},$ (42c)
where $M$ and $L$ are sufficiently small. Combining (5.10) and (5.10) implies that
$|||({\boldsymbol{u}}_{1}-\boldsymbol{u}_{2},{p}_{1}-p_{2},{\theta}_{1}-\theta_{2},{\lambda}_{1}-\lambda_{2})|||^{2}$ $\lesssim 0.$
This completes the proof of the first part. The second part follows from Lemma 37 with the continuity bounds of the bilinear forms and the lifting arguments.

### v2

@@ 4.1.2 The discrete div-conforming problem
(36) $a_{h}(\boldsymbol{u},\boldsymbol{v})$= $\,\mu_{0}\int_{\Omega}\nabla_{h}\boldsymbol{u}:\nabla_{h}\boldsymbol{v}-\sum_{e\in\mathcal{E}_{I}\cup\mathcal{E}_{D}}\mu_{0}\int_{e}\{\!\!\{\nabla_{h}\boldsymbol{u}\}\!\!\}:\llbracket\boldsymbol{v}\otimes\boldsymbol{n}_{e}\rrbracket$
@@
where $\alpha_{0}>0$ is the stabilisation parameter. The broken gradient operator $\nabla_{h}$ is defined by $\nabla_{h}\boldsymbol{u}=\nabla(\boldsymbol{u}|_{K})$ for all $K\in{\mathcal{T}}_{h}$.
For the off-diagonal bilinear form we use the same functional form as $b(\cdot,\cdot)$ but the spaces are different due to the different pairings discussed above
(37) $b_{h}(\boldsymbol{v},(q,\xi)):=-\int_{\Omega}q\nabla\cdot\boldsymbol{v}+{\langle\boldsymbol{v}\cdot\boldsymbol{n},\mathcal{R}_{1/2}\xi\rangle}_{\Gamma_{m}}\qquad\forall\boldsymbol{v}\in\boldsymbol{\mathrm{H}}_{h},(q,\xi)\in Q_{h}\times W_{h}.$
@@
(38) $\widetilde{a}_{h}(\boldsymbol{w};\boldsymbol{u},\boldsymbol{v})=\rho_{0}\int_{\Omega}(\boldsymbol{w}\cdot\nabla_{h}\boldsymbol{u})\cdot\boldsymbol{v}+\frac{\rho_{0}}{2}\sum_{e\in\mathcal{E}_{I}}\int_{e}(\boldsymbol{w}\cdot\boldsymbol{n}_{e}-|\boldsymbol{w}\cdot\boldsymbol{n}_{e}|)(\boldsymbol{u}^{+}-\boldsymbol{u})\cdot\boldsymbol{v},$
where $\boldsymbol{u}^{+}$ is the upwind trace of $\boldsymbol{u}$. If $\boldsymbol{w}\in\boldsymbol{\mathrm{H}}_{h,0}^{0}$, then the following property holds:
@@
(39) $a_{h}(\boldsymbol{u}_{h},\boldsymbol{v}_{h})+\widetilde{a}_{h}(\boldsymbol{u}_{h};\boldsymbol{u}_{h},\boldsymbol{v}_{h})+b_{h}(\boldsymbol{v}_{h},(p_{h},\lambda_{h}))$ $=F(\boldsymbol{v}_{h}),$
@@
(40) $\|\boldsymbol{v}\|_{1,h}^{2}:=\|\boldsymbol{v}\|_{0,\Omega}^{2}+\|\nabla_{h}\boldsymbol{v}\|_{0,\Omega}^{2}+\sum_{e\in\mathcal{E}_{I}\cup\mathcal{E}_{D}}h_{e}^{-1}\|\llbracket\boldsymbol{v}\otimes\boldsymbol{n}_{e}\rrbracket\|^{2}_{0,e}.$
@@
The following bounds hold true
(41a) $|\widetilde{a}(\boldsymbol{w};\boldsymbol{u},\boldsymbol{v})|$ $\lesssim\|\boldsymbol{w}\|_{1,h}\,\|\boldsymbol{u}\|_{1,h}\,\|\boldsymbol{v}\|_{1,h},$ $\boldsymbol{w},\boldsymbol{u}\in\boldsymbol{\mathrm{H}}^{1}(\mathcal{T}_{h}),\boldsymbol{v}\in\boldsymbol{\mathrm{H}}_{h},$
(41b) $|\widetilde{c}(\boldsymbol{w};\theta,\tau)|$ $\lesssim\|\boldsymbol{w}\|_{1,h}\,\|\tau\|_{1,\Omega}\,\,\|\theta\|_{1,\Omega},$ $\boldsymbol{w}\in\boldsymbol{\mathrm{H}}^{1}(\mathcal{T}_{h}),\theta,\tau\in Z_{h},$
@@
$|\widetilde{a}(\boldsymbol{w};\boldsymbol{u},\boldsymbol{v})|\lesssim\|\boldsymbol{w}\|_{\mathbf{L}^{4}(\Omega)}\,\|\boldsymbol{u}\|_{1,h}\,\|\boldsymbol{v}\|_{\mathbf{L}^{4}(\Omega)}\leq\|\boldsymbol{w}\|_{1,h}\,\|\boldsymbol{u}\|_{1,h}\,\|\boldsymbol{v}\|_{1,h}.$
@@ 5 Well-posedness of the divergence-conforming discrete problem
$a_{h}(\boldsymbol{v},\boldsymbol{v})\gtrsim\|\boldsymbol{v}\|_{1,h}^{2}\quad\forall\boldsymbol{v}\in{\boldsymbol{\mathrm{H}}_{h,0}}\quad\mbox{and}\quad c(\tau,\tau)\gtrsim\|\tau\|_{1,\Omega}^{2}\quad\forall\tau\in Z_{h}.$
@@
The first bound directly follows from [19, Prop. 10]. Using (13b) gives the second estimate.
@@
$\sup_{\boldsymbol{v}\in{\boldsymbol{\mathrm{H}}_{h,0}},\boldsymbol{v}\neq\boldsymbol{0}}\frac{b_{h,1}(\boldsymbol{v},q)}{\|\boldsymbol{v}\|_{1,h}}\gtrsim\|q\|_{0,\Omega}\quad\mbox{and}\quad\sup_{\boldsymbol{v}\in{\boldsymbol{\mathrm{H}}_{h,0}},\boldsymbol{v}\neq\boldsymbol{0}}\frac{b_{h,2}(\boldsymbol{v},\xi)}{\|\boldsymbol{v}\|_{1,h}}\gtrsim\|\xi\|_{-1/2,\Gamma_{m,h}},$
for all $q\in Q_{h}$ and for all $\xi\in W_{h}$, where
@@
The first bound directly follows from [19, Prop. 10]. The proof of the second inf-sup condition can be done similarly to [24, Corollary 3.5].
@@
In the following result we prove a global inf-sup condition of the linear part of (39) that shall be useful to ensure the uniqueness and convergence of the discrete solution.
@@
$\lvert\!\lvert\!\lvert(\boldsymbol{v},q,\tau,\xi)\rvert\!\rvert\!\rvert\lesssim\lvert\!\lvert\!\lvert(\boldsymbol{u}_{h},p_{h},\theta_{h},\lambda_{h})\rvert\!\rvert\!\rvert$,
@@
$\lvert\!\lvert\!\lvert(\boldsymbol{u}_{h},p_{h},\theta_{h},\lambda_{h})\rvert\!\rvert\!\rvert^{2}\lesssim B((\boldsymbol{u}_{h},p_{h},\theta_{h},\lambda_{h}),(\boldsymbol{v},q,\tau,\xi)),$
@@
$B(\boldsymbol{u}_{h},p_{h},\theta_{h},\lambda_{h};\boldsymbol{v},q,\tau,\xi):=a_{h}(\boldsymbol{u}_{h},\boldsymbol{v})+b_{h}(\boldsymbol{v},(p_{h},\lambda_{h}))+b_{h}(\boldsymbol{u}_{h},(q,\xi))+c(\theta_{h},\tau),$}
@@
$\lvert\!\lvert\!\lvert(\boldsymbol{v},q,\tau,\xi)\rvert\!\rvert\!\rvert^{2}:=\|\boldsymbol{v}\|_{1,h}^{2}+\|q\|_{0,\Omega}^{2}+\|\xi\|_{-1/2,\Gamma_{m,h}}^{2}+\|\tau\|_{1,\Omega}^{2}.$
@@
Let $(\boldsymbol{u}_{h},p_{h},\theta_{h},\lambda_{h})$ be a solution of (39). Then, the following estimate holds:
$\lvert\!\lvert\!\lvert(\boldsymbol{u}_{h},p_{h},\lambda_{h},\theta_{h})\rvert\!\rvert\!\rvert\lesssim\|\boldsymbol{u}_{in}\|_{1/2,\Gamma_{\mathrm{in}}}+g_{2}.$
Moreover, if $\|\boldsymbol{u}_{h}\|_{1,h}\leq M$, for sufficiently small positive $M<1$, then $(\boldsymbol{u}_{h},p_{h},\theta_{h},\lambda_{h})$ is the unique solution of (39).
@@
The first part follows from Lemma 5.7 with the continuity bounds of the bilinear forms and the lifting arguments. Let $(\boldsymbol{u}_{1},p_{1},\theta_{1},\lambda_{1})$ and $(u_{2},p_{2},\theta_{2},\lambda_{2})$ be two discrete weak solutions of (39). Using Lemma 5.7, for each $({\boldsymbol{u}}_{1}-\boldsymbol{u}_{2},{p}_{1}-p_{2},{\theta}_{1}-\theta_{2},{\lambda}_{1}-\lambda_{2})\in{\boldsymbol{\mathrm{H}}_{h,0}}\times Q_{h}\times Z_{h}\times W_{h}$, we find $(\boldsymbol{v},q,\tau,\xi)\in{\boldsymbol{\mathrm{H}}_{h,0}}\times Q_{h}\times Z_{h}\times W_{h}$ with
$\lvert\!\lvert\!\lvert(\boldsymbol{v},q,\tau,\xi)\rvert\!\rvert\!\rvert\lesssim\lvert\!\lvert\!\lvert({\boldsymbol{u}}_{1}-\boldsymbol{u}_{2},{p}_{1}-p_{2},{\theta}_{1}-\theta_{2},{\lambda}_{1}-\lambda_{2})\rvert\!\rvert\!\rvert$,
@@
$\lvert\!\lvert\!\lvert({\boldsymbol{u}}_{1}-\boldsymbol{u}_{2},{p}_{1}-p_{2},{\theta}_{1}-\theta_{2},{\lambda}_{1}-\lambda_{2})\rvert\!\rvert\!\rvert^{2}\lesssim B({\boldsymbol{u}}_{1}-\boldsymbol{u}_{2},{p}_{1}-p_{2},{\theta}_{1}-\theta_{2},{\lambda}_{1}-\lambda_{2};\boldsymbol{v},q,\tau,\xi).$
By (39), we can assert the bound
$\lvert\!\lvert\!\lvert({\boldsymbol{u}}_{1}-\boldsymbol{u}_{2},{p}_{1}-p_{2},{\theta}_{1}-\theta_{2},{\lambda}_{1}-\lambda_{2})\rvert\!\rvert\!\rvert^{2}$ $\lesssim B({\boldsymbol{u}}_{1},{p}_{1},{\theta}_{1},{\lambda}_{1};\boldsymbol{v},q,\tau,\xi)-B(\boldsymbol{u}_{2},p_{2},\theta_{2},\lambda_{2};\boldsymbol{v},q,\tau,\xi)$
@@
(42) $\quad+|\langle\xi,g(\theta_{1})-g(\theta_{2})\rangle_{\Gamma_{m}}|.$
@@
(43a) $|\widetilde{a}_{h}(\boldsymbol{u}_{1},\boldsymbol{u}_{1},\boldsymbol{v})-\widetilde{a}_{h}(\boldsymbol{u}_{2},\boldsymbol{u}_{2},\boldsymbol{v})|$ $\leq M\|(\boldsymbol{u}_{1}-{\boldsymbol{u}}_{2})\|_{1,h}\;\|\boldsymbol{v}\|_{1,h},$
(43b) $|\widetilde{c}(\boldsymbol{u},\theta,\tau)-\widetilde{c}(\boldsymbol{u},\theta,\tau)|$ $\leq M\big(\|\theta_{1}-{\theta}_{2}\|_{1,\Omega}+\|(\boldsymbol{u}_{1}-{\boldsymbol{u}}_{2})\|_{1,h}\big)\|\tau\|_{1,\Omega},$
(43c) $|\langle\xi,g(\theta_{1})-g(\theta_{2})\rangle_{\Gamma_{m}}|$ $\leq L^{\prime}\|\xi\|_{-1/2,\Gamma_{m,h}}\|(\theta_{1}-{\theta}_{2})\|_{1,\Omega},$
where $M,L^{\prime}>0$ are sufficiently small. Combining (5.10) and (43) implies that
$\lvert\!\lvert\!\lvert({\boldsymbol{u}}_{1}-\boldsymbol{u}_{2},{p}_{1}-p_{2},{\theta}_{1}-\theta_{2},{\lambda}_{1}-\lambda_{2})\rvert\!\rvert\!\rvert^{2}\lesssim 0.$
This completes the proof of the second part.